# Commutation of Smyth and Hoare Power Constructions in Well-filtered Dcpos

### v3

@@ 3.2 The Smyth power construction
Xu et al. proved that for a well-filtered space $X$, $\Sigma\mathcal{Q}(X)$ is still well-filtered [26, Theorem 6.5], which indicates that for each well-filtered dcpo $L$, $\mathcal{Q}(L)$ with the Scott topology is well-filtered, that is, $\mathcal{Q}(L)$ is an object in $\mathbf{WF}$. This fact makes us give a well-defined Smyth power construction.
@@
The Smyth power construction for a well-filtered dcpo is the set of all Scott compact saturated subsets ordered by reverse inclusion.
Recall that, inspired by the facts that for a sober space $X$, the set $\mathcal{Q}(X)$ ordered by reverse inclusion is isomorphic to the set of Scott open filters in $\mathcal{O}(X)$ ordered by inclusion, and for a dcpo $L$, the set of all compact saturated subsets $\mathcal{Q}(L)$ ordered by reverse inclusion is no longer a dcpo, Andrea Schalk focused on the set of all Scott open filters $OFilt(\sigma(L))$, proved that $OFilt(\sigma(L))$ ordered by inclusion remains a dcpo. Concerning this, she gave an endofunctor on the category $\mathbf{DCPO}$ and yields a monad (one can refer to [21, Lemma 7.5, Theorem 7.7]). Fortunately, for a well-filtered dcpo $L$, the set of all Scott compact saturated subsets ordered by reverse inclusion remains a well-filtered dcpo as shown in [26]. Next we will construct an endofunctor $\mathcal{Q}$ on $\mathbf{WF}$ and verify that it is also a monad.
@@
One can easily verify that $\mathcal{Q}$ preserves the identities and the compositions of any two morphisms. So, $\mathcal{Q}$ is an endofunctor on $\mathbf{WF}$.
@@
The saturation is clear. We only consider the compactness. Let $\mathcal{U}$ be a directed subset of $\sigma(L)$ with $\bigcup\mathcal{K}\subseteq\bigcup\mathcal{U}$. Then for each $K\in\mathcal{K}$, its compactness yields the existence of $U_{K}\in\mathcal{U}$ such that $K\subseteq U_{K}$, that is, $K\in\Box U_{K}$. Since the well-filteredness of $L$ implies that each $\Box U$ is Scott open in $\mathcal{Q}(L)$ and $\mathcal{U}$ is directed, $\{\Box U:U\in\mathcal{U}\}$ forms a directed family of Scott open subsets of $\mathcal{Q}(L)$ containing $\mathcal{K}$. So $\mathcal{K}\subseteq\Box U_{0}$ for some $U_{0}\in\mathcal{U}$ by its compactness. Thus $\bigcup\mathcal{K}\subseteq U_{0}$.
@@
The endofunctor $\mathcal{Q}$ with the unit $\theta$ and the multiplication $\iota$ forms a monad on $\mathbf{WF}$, where $\theta$ and $\iota$ are defined as follows when they work on a well-filtered dcpo $L$,
@@
The definitions of $\mathcal{Q}$-algebras reveal that $M$ is a retract of $\mathcal{Q}(M)$. Then by [21, proposition 5.6], we know that $M$ is a deflationary semilattice in $\mathbf{WF}$.
Based on [21, Lemma 4.4] given by Andrea Schalk, we can derive the following results.
@@
Each $\mathcal{Q}$-algebra homomorphism preserves infs of compact sets, hence in particular finite infs.
@@
Let $\mathcal{L}$ be the dcpo given in Section 7.2.4 by Schalk in [21], that is,
@@
$K={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\uparrow}}\{[0,1]\setminus\{x\}:x\in[0,1/2]\}$,
@@ 4 The commutativity between $\mathcal{H}$ and $\mathcal{Q}$
Notably, the forms of these two definitions are similar to those in [4] and [11], where the objects they discussed are quasi-polish spaces and $\mathcal{U}_{S}$-admitting dcpos, respectively. For convenience, we will omit the subscripts when the object acted on is clear.
@@
Then $\phi(\mathcal{A})=\bigcap\{\phi^{\prime}(K):K\in\mathcal{A}\}$ and $\psi(\mathcal{K})=\bigcap\{\psi^{\prime}(A):A\in\mathcal{K}\}$. One can easily check that $\phi^{\prime}(K)=\mathcal{Q}(\eta_{L})(K)={\uparrow}\eta_{L}(K)$ and $\psi^{\prime}(A)=\mathcal{H}(\theta_{L})(A)=cl(\theta_{L}(A))$, which implies that $\phi^{\prime}$ and $\psi^{\prime}$ are well-defined. We know the intersection of arbitrary closed subsets is still closed. So, $\psi$ is well-defined. Now we focus on $\phi$. Lemma 4.2 reveals that $\Sigma\mathcal{H}(L)$ is coherent. Combining with the well-filteredness of $\Sigma\mathcal{H}(L)$, it will be a fact that each intersection of a collection of compact saturated subsets in $\mathcal{H}(L)$ is still compact saturated. Thus $\phi(\mathcal{A})\in\mathcal{Q}(\mathcal{H}(L))$, i.e., $\phi$ is well-defined.
@@
Let $L$ be a well-filtered dcpo. The function $\phi:\mathcal{HQ}(L)\rightarrow\mathcal{QH}(L)$ defined in Definition 4.1 is Scott-continuous.
@@
Assume that there are $\mathcal{A},\mathcal{B}\in\mathcal{HQ}(L)$ with $\mathcal{A}\leq\mathcal{B}$, i.e., $\mathcal{A}\subseteq\mathcal{B}$. By the definition of $\phi$, one can easily see that $\phi(\mathcal{B})\subseteq\phi(\mathcal{A})$, equivalently, $\phi(\mathcal{A})\leq\phi(\mathcal{B})$. So, $\phi$ is order preserving. Let $\{\mathcal{A}_{i}:i\in I\}$ be a directed subset of $\mathcal{HQ}(L)$. To complete this proof, we need to verify that $\phi(\bigvee_{i\in I}\mathcal{A}_{i})=\bigvee_{i\in I}\phi(\mathcal{A}_{i})$, i.e., $\phi(cl(\bigcup_{i\in I}\mathcal{A}_{i}))=\bigcap_{i\in I}\phi(\mathcal{A}_{i})$. Since $\phi$ is order preserving, $\bigcap_{i\in I}\phi(\mathcal{A}_{i})\subseteq\phi(cl(\bigcup_{i\in I}\mathcal{A}_{i}))$ clearly holds. Now pick $A\in\phi(cl(\bigcup_{i\in I}\mathcal{A}_{i}))$, then for each $K\in\bigcup_{i\in I}\mathcal{A}_{i},A\cap K\neq\emptyset$. So we have $A\cap K\neq\emptyset,\forall i\in I,\forall K\in\mathcal{A}_{i}$, which immediately implies that $A\in\phi(\mathcal{A}_{i}),\forall i\in I$. Therefore, $A\in\bigcap_{i\in I}\phi(\mathcal{A}_{i})$ consequently, $\phi(cl(\bigcup_{i\in I}\mathcal{A}_{i}))\subseteq\bigcup_{i\in I}\phi(\mathcal{A}_{i})$ holds.
@@
Recall that a space $X$ is consonant if and only if for any Scott open subset $\mathcal{U}\in\mathcal{O}(X)$ and any $U\in\mathcal{U}$, there exists a $K\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{Q}(X)}$ such that $U\in\Phi(K)\subseteq\mathcal{U}$, where $\Phi(K)=\{V\in\mathcal{O}(X):K\subseteq V\}$. The concept was proposed initially by Dolecki et al. in [5] to answer some questions related to topologies on the hyperspace of closed subsets of a topological space. We say a poset is consonant if it endowed with the Scott topology is consonant.
@@
Let $\mathcal{K}$ be a compact saturated subset of $(\Gamma(P),\sigma(\Gamma(P)))$ and $U$ a Scott open subset of $P$ satisfying that $U\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. Then for each $A\in\mathcal{K}$, $A\in\Diamond U$, which implies that $\mathcal{K}\subseteq\Diamond U$, i.e., $\mathcal{K}\in\Box\Diamond U$. Notably, [4, Theorem 6.13] tells us that the consonance of a topological space $X$ implies the equation $\tau^{-1}(\Diamond\Box U)=\Box\Diamond U$ holds. For the space $\Sigma P$, $\tau$ in [4, Theorem 6.13] is just $\psi$ defined by us. So the consonance of $\Sigma P$ will indicate that $\Box\Diamond U\subseteq\psi^{-1}(\Diamond\Box U)$. Thus $\mathcal{K}\in\psi^{-1}(\Diamond\Box U)$. From which we have that $\psi(\mathcal{K})\in\Diamond\Box U$, i.e., $\psi(\mathcal{K})\cap\Box U\neq\emptyset$. So there exists a $K\in\psi(\mathcal{K})$ contained in $U$. Therefore, we find a compact saturated subset $K$ contained in $U$ that meets each element of $\mathcal{K}$. Hence, $P$ has the property $(\mathrm{KC})$.
Now see [4, Theorem 6.10] again, it states that for a space $X$, $\sigma(\mathcal{O}(X))$ and $\mathcal{Q}(\Gamma(X))$ are isomorphic lattices. Recently, Chen et al. in [3] proved that if a poset $P$ under the Scott topology is core-compact, then $\sigma(\Gamma(P))=\nu(\Gamma(P))$. Thus for a poset $P$, $\sigma(\sigma(P))$ would be isomorphic to $\mathcal{Q}(\Gamma(P))$ as long as $\Sigma P$ is core-compact. Combining all of these results and the above lemma, we obtain the following result.
@@
$(\Rightarrow)$: Let $\mathcal{U}$ be a Scott open subset of $\mathcal{Q}(L)$ and $K$ an element in $\mathcal{U}$. Then $K$ does not belong to the Scott closed subset $\mathcal{Q}(L)\setminus\mathcal{U}$. Since $\psi$ is surjective, there exists a $\mathcal{K}\in\mathcal{Q}(\mathcal{H}(L))$ such that $\mathcal{Q}(L)\setminus\mathcal{U}=\psi(\mathcal{K})$, then $K\notin\psi(\mathcal{K})$. By the definition of $\psi$, it yields that $K\cap A=\emptyset$ for some $A\in\mathcal{K}$. The Scott open subset $V=L\setminus A$ contains $K$, that is, $K\in\Box V$. Thus we only need to show $\Box V\subseteq\mathcal{U}$. For any $M\in\Box V$, $M\cap A=\emptyset$, which implies that $M\notin\psi(\mathcal{K})$. Hence, $M\in\mathcal{U}$.
Let $\mathcal{K}$ be a Scott compact saturated subset of $\Gamma(L)$ and $U\in\sigma(L)$ satisfying $U\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. Because of $U\cap(L\setminus U)=\emptyset$, we have $L\setminus U\notin\mathcal{K}$. The fact that $\phi$ is surjective suggests that there is an $\mathcal{A}\in\mathcal{H}(\mathcal{Q}(L))$ such that $\mathcal{K}=\phi(\mathcal{A})$. Thus $L\setminus U\notin\phi(\mathcal{A})$. This means $(L\setminus U)\cap K_{0}=\emptyset$ i.e., $K_{0}\subseteq U$ for some $K_{0}\in\mathcal{A}$. Now suppose that $K_{0}\cap A_{1}=\emptyset$ for some $A_{1}\in\mathcal{K}$. Then $A_{1}\notin\phi(\mathcal{A})=\mathcal{K}$, which is an contradiction obviously. Thus $K_{0}$ intersects each $A\in\mathcal{K}$ and $L$ has the property $(\mathrm{KC})$ as desired.
$(\Leftarrow)$: Obviously, $\phi$ and $\psi$ are order-preserving. Since order-preserving isomorphisms between dcpos are Scott-continuous, the proof will be complete after showing that $\psi\circ\phi$ and $\phi\circ\psi$ are identities, more precisely, the following two equations hold (for arbitrary $\mathcal{A}\in\mathcal{H}(\mathcal{Q}(L))$ and $\mathcal{K}\in\mathcal{Q}(\mathcal{H}(L))$):
@@
Note that the definitions of $\phi$ and $\psi$ determine that $\mathcal{A}\subseteq\psi\circ\phi(\mathcal{A})$ and $\mathcal{K}\subseteq\phi\circ\psi(\mathcal{K})$. Thus we just need to check the reverse.
As to the first one, suppose that there is a $K\in\mathcal{Q}(L)$ belonging to $\psi\circ\phi(\mathcal{A})$ but not to $\mathcal{A}$. Then $K\in\mathcal{Q}(L)\setminus\mathcal{A}$, which is Scott open in $\mathcal{Q}(L)$ and further an open subset of $\mathcal{Q}(L)$ with respect to the upper Vietoris topology by $(\mathrm{i})$. Thus there exists a $U\in\sigma(L)$ such that $K\in\Box U\subseteq\mathcal{Q}(L)\setminus\mathcal{A}$, which implies $K\cap(L\setminus U)=\emptyset$. This will contradict the fact that $K\in\psi\circ\phi(\mathcal{A})$ if we could show $L\setminus U\in\phi(\mathcal{A})$. Assume $(L\setminus U)\cap K_{0}=\emptyset$ for some $K_{0}\in\mathcal{A}$. Then $K_{0}\in\Box U$, thus $K_{0}\notin\mathcal{A}$, a contradiction. So $L\setminus U\in\phi(\mathcal{A})$. Therefore, $\psi\circ\phi(\mathcal{A})\subseteq\mathcal{A}$.
Regarding the second, for the sake of a contradiction, we assume that there is an $A\in\phi\circ\psi(\mathcal{K})$ but $A\notin\mathcal{K}$. It suggests that each $C\in\mathcal{K}$ is not contained in $A$, that is to say, $(L\setminus A)\cap C\neq\emptyset$ for all $C\in\mathcal{K}$. Since $L$ has the property $(\mathrm{KC})$, we can find a $K_{0}\subseteq L\setminus A$ intersecting each $C$ in $\mathcal{K}$. Then $K_{0}\in\psi(\mathcal{K})$ and $K_{0}\cap A=\emptyset$, which contradicts that $A\in\phi\circ\psi(\mathcal{K})$. Thus $A\in\mathcal{K}$, which obviously leads to a contradiction. So our assumption is wrong and $\phi\circ\psi(\mathcal{K})\subseteq\mathcal{K}$ holds.
It was Schalk who proved that for a locally compact sober space $X$, the upper Vietoris topology and the Scott topology on $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{Q}(X)$} coincide (see [21, Lemma 7.26]), meanwhile, [3, Proposition 5.4] reveals that each locally compact space is consonant. Therefore, combining Lemma 4.6 and Theorem 4.8, we further draw the conclusion in the following.

### v5

@@ 3.2 The Smyth power construction
Xu et al. proved that for a well-filtered space $X$, $\Sigma\mathcal{Q}(X)$ remains well-filtered [26, Theorem 6.5], which indicates that for each well-filtered dcpo $L$, $\mathcal{Q}(L)$ with the Scott topology is well-filtered, that is, $\mathcal{Q}(L)$ is an object in $\mathbf{WF}$. This fact gives us a well-defined Smyth power construction.
@@
The Smyth power construction for a well-filtered dcpo $L$ is the set $\mathcal{Q}(L)$ of all Scott compact saturated subsets ordered by reverse inclusion.
Recall that, inspired by the facts that for a sober space $X$, the set $\mathcal{Q}(X)$ ordered by reverse inclusion is isomorphic to the set of Scott open filters in $\mathcal{O}(X)$ ordered by inclusion, and for a general dcpo $L$, the set $\mathcal{Q}(L)$ of all compact saturated subsets ordered by reverse inclusion is no longer a dcpo, A. Schalk focused on the set of all Scott open filters $OFilt(\sigma(L))$ and proved that $OFilt(\sigma(L))$ ordered by inclusion remains a dcpo. Concerning this, she gave an endofunctor on the category $\mathbf{DCPO}$ and obtained a monad (one can refer to [21, Lemma 7.5, Theorem 7.7]). Fortunately, for a well-filtered dcpo $L$, the set of all Scott compact saturated subsets ordered by reverse inclusion remains a well-filtered dcpo as shown in [26]. Next we will construct an endofunctor $\mathcal{Q}$ on $\mathbf{WF}$ and verify that it is also a monad.
@@
One can easily verify that $\mathcal{Q}$ preserves the identities and the compositions of any two morphisms. So $\mathcal{Q}$ is an endofunctor on $\mathbf{WF}$.
@@
The saturation is clear, so we only need to consider the compactness. Let $\mathcal{U}$ be a directed subset of $\sigma(L)$ with $\bigcup\mathcal{K}\subseteq\bigcup\mathcal{U}$. Then for each $K\in\mathcal{K}$, its compactness yields the existence of $U_{K}\in\mathcal{U}$ such that $K\subseteq U_{K}$, that is, $K\in\Box U_{K}$. Since the well-filteredness of $L$ implies that each $\Box U$ is Scott open in $\mathcal{Q}(L)$ and $\mathcal{U}$ is directed, $\{\Box U:U\in\mathcal{U}\}$ forms a directed family of Scott open subsets of $\mathcal{Q}(L)$, and $\mathcal{K}\subseteq\bigcup\{\Box U:U\in\mathcal{U}\}$. As $\mathcal{K}$ is compact, there exists a $U_{0}\in\mathcal{U}$ such that $\mathcal{K}\subseteq\Box U_{0}$, which immediately implies $\bigcup\mathcal{K}\subseteq U_{0}$. So $\bigcup\mathcal{K}$ is compact.
@@
The endofunctor $\mathcal{Q}$ with the unit $\theta:Id\rightarrow\mathcal{Q}$ and the multiplication $\iota:\mathcal{Q}^{2}\rightarrow\mathcal{Q}$ forms a monad on $\mathbf{WF}$, where $\theta$ and $\iota$ are defined as follows when they act on a well-filtered dcpo $L$:
@@
The definition of $\mathcal{Q}$-algebras reveals that $M$ is a retract of $\mathcal{Q}(M)$. Then by [21, Proposition 5.6], we know that $M$ is a deflationary semilattice in $\mathbf{WF}$.
Based on [21, Lemma 4.4] given by A. Schalk, we can derive the following results.
@@
Each $\mathcal{Q}$-algebra homomorphism preserves infs of compact sets; hence, in particular finite infs.
@@
Let $\mathcal{L}$ be the dcpo given in Section 7.2.4 by A. Schalk in [21], that is,
@@
$K={\uparrow}\{[0,1]\setminus\{x\}:x\in[0,1/2]\}$,
@@ 4 The commutativity between $\mathcal{H}$ and $\mathcal{Q}$
Notably, these two definitions are similar in form to those given in [4] and [11], where the authors give the definitions and discuss them for all topological spaces and $\mathcal{U}_{S}$-admitting dcpos, respectively. For convenience, we will omit the subscripts when the object acted on is clear.
@@
Then $\phi(\mathcal{A})=\bigcap\{\phi^{\prime}(K):K\in\mathcal{A}\}$ and $\psi(\mathcal{K})=\bigcap\{\psi^{\prime}(A):A\in\mathcal{K}\}$. One can easily check that $\phi^{\prime}(K)=\mathcal{Q}\eta_{L}(K)={\uparrow}\eta_{L}(K)$ and $\psi^{\prime}(A)=\mathcal{H}\theta_{L}(A)=cl(\theta_{L}(A))$, which implies that $\phi^{\prime}$ and $\psi^{\prime}$ are well-defined. We know the intersection of arbitrary closed subsets is still closed. So $\psi$ is well-defined. Now we focus on $\phi$. Lemma 4.2 reveals that $\Sigma\mathcal{H}(L)$ is coherent. Combining with the well-filteredness of $\Sigma\mathcal{H}(L)$, one can obtain that each intersection of a collection of compact saturated subsets in $\mathcal{H}(L)$ is still compact saturated. Thus $\phi(\mathcal{A})\in\mathcal{Q}(\mathcal{H}(L))$, i.e., $\phi$ is well-defined.
@@
Let $L$ be a well-filtered dcpo. The function $\phi:\mathcal{HQ}(L)\rightarrow\mathcal{QH}(L)$ defined in Definition 4.1 preserves all sups, hence is Scott-continuous.
@@
Assume that there are $\mathcal{A},\mathcal{B}\in\mathcal{HQ}(L)$ with $\mathcal{A}\leq\mathcal{B}$, i.e., $\mathcal{A}\subseteq\mathcal{B}$. By the definition of $\phi$, one can easily see that $\phi(\mathcal{B})\subseteq\phi(\mathcal{A})$, equivalently, $\phi(\mathcal{A})\leq\phi(\mathcal{B})$. So $\phi$ is order-preserving. Let $\{\mathcal{A}_{i}:i\in I\}$ be an arbitrary subset of $\mathcal{HQ}(L)$. To complete this proof, we need to verify that $\phi(\bigvee_{i\in I}\mathcal{A}_{i})=\bigvee_{i\in I}\phi(\mathcal{A}_{i})$. Since $\phi$ is order-preserving, for each $i\in I$, $\phi(\mathcal{A}_{i})\leq\phi(\bigvee_{i\in I}\mathcal{A}_{i})$. Then $\bigvee_{i\in I}\phi(\mathcal{A}_{i})\leq\phi(\bigvee_{i\in I}\mathcal{A}_{i})$ holds. Now we consider the reverse, more precisely, to prove $\bigcap_{i\in I}\phi(\mathcal{A}_{i})\subseteq\phi(cl(\bigcup_{i\in I}\mathcal{A}_{i}))$. Take any closed subset $A\in\bigcap_{i\in I}\phi(\mathcal{A}_{i})$. Suppose that $A\notin\phi(cl(\bigcup_{i\in I}\mathcal{A}_{i}))$, i.e., there exists a compact saturated subset $K_{0}\in cl(\bigcup_{i\in I}\mathcal{A}_{i})$ such that $A\cap K_{0}=\emptyset$. Then $K_{0}\in\Box(L\setminus A)$, where $\Box(L\setminus A)$ is a Scott open subset of $\mathcal{Q}(L)$ by the well-filteredness of $L$. So $\Box(L\setminus A)\bigcap(\bigcup_{i\in I}\mathcal{A}_{i})\neq\emptyset$, which implies that there is an $i_{0}\in I$ such that $\Box(L\setminus A)\bigcap\mathcal{A}_{i_{0}}\neq\emptyset$. That is to say, we can find a $K_{1}\in\mathcal{A}_{i_{0}}$ contained in $L\setminus A$. This immediately indicates that $K_{1}\cap A=\emptyset$; thus $A\notin\phi(\mathcal{A}_{i_{0}})$, which contracts $A\in\bigcap_{i\in I}\phi(\mathcal{A}_{i})$. Therefore, $A\in\phi(cl(\bigcup_{i\in I}\mathcal{A}_{i}))$. From the arbitrariness of $A$, we conclude that $\bigcap_{i\in I}\phi(\mathcal{A}_{i})\subseteq\phi(cl(\bigcup_{i\in I}\mathcal{A}_{i}))$, that is, $\phi(\bigvee_{i\in I}\mathcal{A}_{i})\leq\bigvee_{i\in I}\phi(\mathcal{A}_{i})$ holds.
@@
Recall that a space $X$ is consonant if and only if for any Scott open subset $\mathcal{U}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\subseteq}\mathcal{O}(X)$ and any $U\in\mathcal{U}$, there exists a $K\in\mathcal{Q}(X)$ such that $U\in\Phi(K)\subseteq\mathcal{U}$, where $\Phi(K)=\{V\in\mathcal{O}(X):K\subseteq V\}$. The concept was proposed initially by Dolecki et al. in [5] to answer some questions related to topologies on the hyperspace of closed subsets of a topological space. We say a poset is consonant if it endowed with the Scott topology is consonant.
@@
Let $\mathcal{K}$ be a compact saturated subset of $(\Gamma(P),\sigma(\Gamma(P)))$ and $U$ a Scott open subset of $P$ satisfying $U\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. Set $\mathcal{U}=\bigcap\{\vartriangle\!\!A:A\in\mathcal{K}\}$, where $\vartriangle\!\!A=\{V\in\sigma(P):V\cap A\neq\emptyset\}$. We claim that $\mathcal{U}\in\sigma(\sigma(P))$. Obviously, it is an upper set. Assume $\mathcal{V}$ is a directed subset of $\sigma(P)$ and $\sup\mathcal{V}=\bigcup\mathcal{V}\in\mathcal{U}$. Then $\bigcup\mathcal{V}\in\vartriangle\!\!A$ for all $A\in\mathcal{K}$, that is, $\bigcup\mathcal{V}\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. It means that for each $A\in\mathcal{K}$, there exists a $V_{A}\in\mathcal{V}$ such that $V_{A}\cap A\neq\emptyset$, i.e., $A\in\lozenge V_{A}$, and thus $\mathcal{K}\subseteq\bigcup\{\lozenge V:V\in\mathcal{V}\}$. By the compactness of $\mathcal{K}$ and the directedness of $\{\lozenge V:V\in\mathcal{V}\}$, there must exist a $V_{0}\in\mathcal{V}$ such that $\mathcal{K}\subseteq\lozenge V_{0}$. It follows that $V_{0}\in\vartriangle\!\!A$ for each $A\in\mathcal{K}$, that is to say, $V_{0}\in\mathcal{U}$. Thus $\mathcal{U}$ is Scott open. Since $P$ is consonant and $U\in\mathcal{U}$, there exists a $K\in\mathcal{Q}(P)$ satisfying $U\in\Phi(K)\subseteq\mathcal{U}$. We will complete the proof if we show that $K\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. Suppose $K\cap A_{0}=\emptyset$ for some $A_{0}\in\mathcal{K}$. Then $K\subseteq P\setminus A_{0}$, where $P\setminus A_{0}\in\sigma(P)$. This implies $P\setminus A_{0}\in\Phi(K)\subseteq\mathcal{U}$. So $P\setminus A_{0}\in\;\vartriangle\!\!A_{0}$, which contradicts the definition of $\vartriangle\!\!A_{0}$.
Note that the property $(\mathrm{KC})$ for $T_{0}$ spaces is exactly the condition $\mathcal{K}\in\Box\diamondsuit U$ implies $\tau(\mathcal{K})\in\diamondsuit\Box U$” (i.e., $\Box\diamondsuit U\subseteq\tau^{-1}(\diamondsuit\Box U)$ in the notation of [4], where the $\tau$ in [4] is the same as the $\psi$ defined in Definition 4.1. Since Lemma 6.5 in [4] has shown that $\tau^{-1}(\diamondsuit\Box U)\subseteq\Box\diamondsuit U$, the following lemma is equivalent to the fact that (1) $\Leftrightarrow$ (3) in [4, Theorem 6.13].
Lemma 4.7.
A $T_{0}$ space $X$ is consonant if and only if it has the property $(\mathrm{KC})$.
Now see [4, Theorem 6.10], it states that for a space $X$, $\sigma(\mathcal{O}(X))$ and $\mathcal{Q}(\Gamma(X))$ are isomorphic lattices. Recently, Chen et al. in [3] proved that if a poset $P$ endowed with the Scott topology is core-compact, then $\sigma(\Gamma(P))=\nu(\Gamma(P))$. Thus for a poset $P$, $\sigma(\sigma(P))$ would be isomorphic to $\mathcal{Q}(\Gamma(P))$ as long as $\Sigma P$ is core-compact. Combining all of these results and the above lemma, we obtain the following result.
@@
$(\Rightarrow)$: (i) Let $\mathcal{U}$ be a Scott open subset of $\mathcal{Q}(L)$ and $K$ an element in $\mathcal{U}$. Then $K\notin\mathcal{Q}(L)\setminus\mathcal{U}$, where $\mathcal{Q}(L)\setminus\mathcal{U}$ is a closed subset of $\Sigma\mathcal{Q}(L)$. Since $\psi$ is surjective, there exists a $\mathcal{K}\in\mathcal{Q}(\mathcal{H}(L))$ such that $\mathcal{Q}(L)\setminus\mathcal{U}=\psi(\mathcal{K})$; consequently, $K\notin\psi(\mathcal{K})$. From the definition of $\psi$, we have $K\cap A=\emptyset$ for some $A\in\mathcal{K}$. Set $V=L\setminus A$. Then $V\in\sigma(L)$ and $K\subseteq V$, i.e., $K\in\Box V$. We now show $\Box V\subseteq\mathcal{U}$. Take any $M\in\Box V$. Then $M\cap A=\emptyset$, which implies $M\notin\psi(\mathcal{K})=\mathcal{Q}(L)\setminus\mathcal{U}$; hence, $M\in\mathcal{U}$. Therefore $K\in\Box V\subseteq\mathcal{U}$, so $\mathcal{U}\in\upsilon(Q(L))$. By the arbitrariness of $\mathcal{U}$, we conclude that $\sigma(Q(L))\subseteq\upsilon(Q(L))$.
(ii) Let $\mathcal{K}$ be a Scott compact saturated subset of $\Gamma(L)$ and $U\in\sigma(L)$ with $U\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. Since $U\cap(L\setminus U)=\emptyset$, we have $L\setminus U\notin\mathcal{K}$. By the surjectivity of $\phi$, there exists an $\mathcal{A}\in\mathcal{H}(\mathcal{Q}(L))$ such that $\mathcal{K}=\phi(\mathcal{A})$. Thus $L\setminus U\notin\phi(\mathcal{A})$. This means $(L\setminus U)\cap K_{0}=\emptyset$, i.e., $K_{0}\subseteq U$ for some $K_{0}\in\mathcal{A}$. Now suppose that $K_{0}\cap A_{1}=\emptyset$ for some $A_{1}\in\mathcal{K}$. Then $A_{1}\notin\phi(\mathcal{A})=\mathcal{K}$, which is a contradiction obviously. Therefore, $K_{0}$ intersects every element in $\mathcal{K}$, and $L$ has the property $(\mathrm{KC})$ as desired.
$(\Leftarrow)$: Clearly, $\phi$ and $\psi$ are order-preserving. Since order-preserving isomorphisms between dcpos are Scott-continuous, the proof will be complete once we show that $\psi\circ\phi$ and $\phi\circ\psi$ are identity maps, more precisely, the following two equations hold for arbitrary $\mathcal{A}\in\mathcal{H}(\mathcal{Q}(L))$ and $\mathcal{K}\in\mathcal{Q}(\mathcal{H}(L))$:
@@
It follows directly from the definitions of $\phi$ and $\psi$ that $\mathcal{A}\subseteq\psi\circ\phi(\mathcal{A})$ and $\mathcal{K}\subseteq\phi\circ\psi(\mathcal{K})$. Thus we just need to check the reverse inclusions.
For the first inclusion, suppose that there is a compact saturated subset $K$ belonging to $\psi\circ\phi(\mathcal{A})$ but not to $\mathcal{A}$. Then $K\in\mathcal{Q}(L)\setminus\mathcal{A}$, which is Scott open in $\mathcal{Q}(L)$ and by $(\mathrm{i})$, is also open in the upper Vietoris topology. Hence, there exists a $U\in\sigma(L)$ such that $K\in\Box U\subseteq\mathcal{Q}(L)\setminus\mathcal{A}$, which implies $K\cap(L\setminus U)=\emptyset$. This would contradict the fact that $K\in\psi\circ\phi(\mathcal{A})$ once we show $L\setminus U\in\phi(\mathcal{A})$. Assume $(L\setminus U)\cap K_{0}=\emptyset$ for some $K_{0}\in\mathcal{A}$. Then $K_{0}\in\Box U$, from which we have $K_{0}\notin\mathcal{A}$, a contradiction. Thus $L\setminus U\in\phi(\mathcal{A})$; consequently, $\psi\circ\phi(\mathcal{A})\subseteq\mathcal{A}$ holds.
To prove the second inclusion, for the sake of a contradiction, we assume that there is an $A\in\phi\circ\psi(\mathcal{K})$ but $A\notin\mathcal{K}$. This indicates that there is no $C\in\mathcal{K}$ contained in $A$, that is, $(L\setminus A)\cap C\neq\emptyset$ for all $C\in\mathcal{K}$. Since $L$ has the property $(\mathrm{KC})$, we can find a $K_{0}\subseteq L\setminus A$ that intersects each $C$ in $\mathcal{K}$. Thus $K_{0}\in\psi(\mathcal{K})$ and $K_{0}\cap A=\emptyset$, which contradicts that $A\in\phi\circ\psi(\mathcal{K})$. So each element in $\phi\circ\psi(\mathcal{K})$ belongs to $\mathcal{K}$, that is, $\phi\circ\psi(\mathcal{K})\subseteq\mathcal{K}$.
It was A. Schalk who proved that for a locally compact sober space $X$, the upper Vietoris topology and the Scott topology on $\mathcal{Q}(X)$ coincide (see [21, Lemma 7.26]), meanwhile, [3, Proposition 5.4] reveals that each locally compact space is consonant. Therefore, combining Lemma 4.6 and Theorem 4.9, we further draw the conclusion in the following.